OpenStudy (anonymous):

2/3y+6=y a)-18 b)18 c)1/3 d)-1/3

OpenStudy (anonymous):

@terenzreignz @Ashley1997

terenzreignz (terenzreignz):

(2/3y) is a fraction?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, try multiplying 3y to both sides of the equation, then

OpenStudy (anonymous):

B?

terenzreignz (terenzreignz):

Or, there are only four choices after all, substitute them one by one and see which one works :P

terenzreignz (terenzreignz):

But I don't think any of them really works...

OpenStudy (anonymous):

Well one of them is correct....so..

OpenStudy (anonymous):

Whatd you get @giovkast

OpenStudy (anonymous):

2/3y + 6 = y 6 = y - (2/3)y 6 = (1/3)y then you multiply both sides by 3 18 = y

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

yes, thank u @giovkast

terenzreignz (terenzreignz):

but I thought it was (2/3y) and not (2/3)y... derr sorry :)

OpenStudy (anonymous):

np